OpenStudy (anonymous):

Proofing: sinxcosx/1-2cos^2x = 1/tanx-cotx Help :c

OpenStudy (aum):

\[ \frac{\sin x \cos x}{1-2\cos^2 x} = \frac{\sin x \cos x}{\sin^2 x + \cos^2 x-2\cos^2 x} = \frac{\sin x \cos x}{\sin^2 x - \cos^2 x} = \\ \text{ } \\ \Large \frac{\frac {\sin x \cos x}{\sin x \cos x}}{\frac{\sin^2 x - \cos^2 x}{\sin x \cos x}} = \frac{1}{\tan x - \cot x} \]

OpenStudy (anonymous):

Thanks so much! I understand now ^^

OpenStudy (aum):

You are welcome.